# Confidence-aware Self-Semantic Distillation on Knowledge Graph Embedding

### v2

@@ Abstract.
Knowledge Graph Embedding (KGE), which projects entities and relations into continuous vector spaces, have garnered significant attention. Although high-dimensional KGE methods offer better performance, they come at the expense of significant computation and memory overheads. Decreasing embedding dimensions significantly deteriorates model performance. While several recent efforts utilize knowledge distillation or non-Euclidean representation learning to augment the effectiveness of low-dimensional KGE, they either necessitate a pre-trained high-dimensional teacher model or involve complex non-Euclidean operations, thereby incurring considerable additional computational costs. To address this, this work proposes Confidence-aware Self-Knowledge Distillation (CSD) that learns from model itself to enhance KGE in a low-dimensional space. Specifically, CSD extracts knowledge from embeddings in previous iterations, which would be utilized to supervise the learning of the model in the next iterations. Moreover, a specific semantic module is developed to filter reliable knowledge by estimating the confidence of previously learned embeddings. This straightforward strategy bypasses the need for time-consuming pre-training of teacher models and can be integrated into various KGE methods to improve their performance. Our comprehensive experiments on six KGE backbones and four datasets underscore the effectiveness of the proposed CSD.
Knowledge Graph Embedding, Knowledge Distillation, Self-distillation, Confidence
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Data mining
@@ 1. Introduction
Knowledge Graphs (KGs), which encapsulate real-world facts in a graph structure, have found extensive applications across various domains, ranging from event prediction (EP1, ; EP2, ), question answering (QE1, ; QE2, ), to recommender systems (RS1, ; RS2, ). As a foundational task in this area, Knowledge Graph Embedding (KGE) (TransE, ) aims to represent entities and relations in KGs as continuous vectors. With the increasing scale of KGs in practice, the necessity for both accuracy and efficiency in KGE becomes increasingly critical. While recent years have witnessed a proliferation of sophisticated KGE methods, most successes hinge on high-dimensional embeddings, which inevitably incur increased memory and computational overhead. For instance, as demonstrated in Table 1, representative KGE methods such as LowFER (LowFER, ) and DistMult (DistMult, ) exhibit a significant performance decline (96%) in their low-dimensional versions (e.g., 25 dimensions) compared to their high-dimensional (e.g., 250 dimensions) counterparts, despite the low-dimensional models have 6-11 times fewer embedding layer parameters and offering 3-5 times improved inference latency.
To enhance the effectiveness of low-dimensional KGE, several recent efforts have emerged, which can be categorized into two main approaches: 1) leveraging knowledge distillation to empower a low-dimensional model with knowledge from a high-dimensional teacher model (liu2023iterde, ; MulDE, ; DualDE, ); and 2) embedding KGs in non-Euclidean spaces (e.g., hyperbolic or mixed-curvature spaces) to better capture the complex patterns of KGs (cao2022geometry, ). However, these strategies still face limitations: knowledge distillation requires pre-training a high-dimensional KGE model, which still demands extensive computational resources and memory; non-Euclidean methods involve complex operations that incur additional computational costs and lack flexibibity to be applied in various KGE models (yao2023analogical, ). Given these limitations, an intriguing research question arises: How can we achieve effective low-dimensional KGE without incurring significant additional training costs?
@@
To this end, this work proposes to leverage self-knowledge distillation (SKD) for both efficient and effective low-dimensional knowledge graph embedding. Unlike traditional knowledge distillation (KD), which requires a pre-trained teacher model, SKD learns from the model itself through introspection (selfKD1, ; selfKD2, ; snapshot, ; yun2020regularizing, ). Specifically, SKD utilizes information from past epochs to extract valuable knowledge to guide learning in subsequent epochs. SKD has been proven to enhance model accuracy and robustness while avoiding the substantial training overhead, and has been successfully applied in fields such as computer vision and natural language processing. However, adapting sophisticated SKD strategies to KGE presents unique challenges:
@@
  • We propose a novel iterative self-knowledge distillation method tailored for knowledge graph embedding that employs a specific semantic extraction module to distill reliable embedding knowledge from previous epochs to guide model learning.
@@ 2.1. Knowledge Graph Embedding
Recently, there has been intensive research on a neural network model called ConvE (ConvE, ), which aims to capture the semantic interactions between entities and relations in knowledge graphs. Building upon this, ConvR (ConvR, ) improves ConvE by replacing global convolution filters with relation-based filters. These filters are adaptively generated from relation embeddings, allowing ConvR to capture entity-relation interactions in knowledge graphs. Another model called HypER (HypER, ) further extends ConvE by using a hypernetwork to generate relation-specific convolution filters through tensor factorization. This extension facilitates feature interactions between entities and relations in knowledge graphs. Additionally, AcrE (AcrE, ) and InteractE (InteractE, ) introduce various convolution mechanisms, such as atrous and circular convolution, to enhance entity and relation semantic interactions. JointE (JointE, ) takes this further by jointly utilizing 1D and 2D convolutions to capture translation properties in KGs, thereby increasing interactions between entities and relations. M-DCN (M-DCN, ) introduces a dynamic and adaptive generation of multi-scale filters for convolutions. It captures feature interactions and improves model expressiveness in knowledge graphs.
@@ 2.2. Knowledge Distillation
To get both of performance and inference speed, knowledge distillation technology has been proposed. The initial knowledge distillation approach focused on utilizing the teacher model’s prediction probabilities as soft targets for training the student model (firstKD, ). FitNet (FitNet, ) took a step further by incorporating the teacher model’s intermediate features as hints for training the student model. Building on FitNet, SemCKD (SemCKD, ) introduced semantic calibration, which employed an attention mechanism to obtain adaptive intermediate features for student training. Additionally, SimKD (SimKD, ) achieved competitive performance for the student model simply by reusing the discriminant classifier from the teacher model.
The pioneer in knowledge graph embedding, KD-MKB (KD-MKB, ), first harnessed knowledge distillation for knowledge graph embedding by establishing a cooperative learning framework between a teacher model and a student model. Following this, MulDE (MulDE, ) and DualDE (DualDE, ) are proposed. All these papers utilized high-dimensional, powerful teacher models to facilitate training of student models for improved performance. However, a high-capacity teacher model is not available in many cases. To address this limitation, we shifting our focus towards the self-knowledge distillation strategy(selfKD1, ; selfKD2, ; furlanello2018born, ; yun2020regularizing, ). However, existing self-distillation methods predominantly cater to the field of computer vision and perform sub-optimally within the field of knowledge graph embedding, where they manage input data that generally assume independence from each other. However, in the Knowledge Graph Embedding (KGE) domain, the data that we input are interconnected. The current self-distillation algorithms fail to leverage this aspect of information. Consequently, to harness this component of data, we propose the Improved Self-Distillation (CSD) method.
@@ 3.1. Knowledge Graph Embedding
A knowledge graph (YAGO, ), denoted as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{T}\}$, encompasses substantial quantities of entities and relations. Herein, $\mathcal{V}$, $\mathcal{R}$ and $\mathcal{T}$ signify the sets of entities, relations and triplets respectively. Each fact $t\in\mathcal{T}$, is presented as a triple $(h,r,t)\in\mathcal{V}\times\mathcal{R}\times\mathcal{V}$, describing there is a relationship $r\in\mathcal{R}$ from the head entity $h\in\mathcal{V}$ to the tail entity $t\in\mathcal{V}$. The target of Knowledge graph embedding (KGE) is to project each entity $v\in\mathcal{V}$ and relation $r\in\mathcal{R}$ into corresponding d-dimensional representations $\mathbf{e}_{v},\mathbf{f}_{r}$. For convenient, we collect the embeddings of all entities and relations as matrices $\mathbf{E}$ and $\mathbf{F}$, respectively. These embeddings are subsequently utilized to facilitate downstream tasks.
To derive embeddings from Knowledge Graphs (KGs), existing Knowledge Graph Embedding (KGE) methods (TuckER, ; DistMult, ; ComplEx, ) typically formulate the task as a specific link prediction problem. These methods establish a scoring function $S_{hrt}=g(\mathbf{e}_{h},\mathbf{f}_{r},\mathbf{e}_{t})$ based on embeddings to estimate the validity of a triplet. For instance, TransE (TransE, ) adopts $g(\mathbf{e}_{h},\mathbf{f}_{r},\mathbf{e}_{t})=-\|\mathbf{e}_{h}+\mathbf{f}_{r}-\mathbf{e}_{t}\|_{1/2}$. Typically, the BCE loss can be used to optimize the KGE models as follows:
@@ 3.2. Self-Knowledge Distillation
(3) $D_{H}(a,b)=\begin{cases}\frac{1}{2}*(a-b)^{2},&|a-b|\leq 1\\ |\mathbf{a}-b|-\frac{1}{2},&|a-b|>1.\end{cases}$
SKD employs a previous snapshot of the model as the teacher to supervise the learning of the current model.
@@ 4. Methodology
In this section, we detail the proposed Confidence-aware Self-Semantic Distillation (CSD), which is tailored for low-dimensional KGE. As shown in Fig. 2 (a), CSD follows conventional architecture of SKD, where a KGE model alternates between the roles of teacher and student during the training process. The knowledge of the previous iteration would be utilized to supervise the learning the current model. However, given the specific challenges in KGE, we make specific designs to adapt SKD to this task.
Diverge from utilizing soft labels in the classification task, we distill knowledge from the model embeddings, by taking model status in $l-1$-th iteration as a teacher in $l$-th iteration. A specific semantic extraction block is developed to extract reliable knowledge from model embeddings to supervise the model learning. Semantic information is dynamically and adaptively generated by the semantic extraction block, showcasing dynamic and evolution properties throughout the training process (discussed in Section 4.1). Throughout these iterations, we leverage the teacher model to transfer semantic information to the student model by optimizing the Huber Loss between their semantic information. This optimization facilitates the student model to progressively incorporate and accumulate semantic knowledge from the past model.
@@ 4.1. Semantic Extraction Block
The embedding matrices $E$ and $F$ encapsulate the knowledge graph’s semantic attributes based on the entities and relations. However, we find that directly distilling the embedding matrices between the teacher and student models leads to poor performance gain. This is due to the complex structures and inherently noisy nature of KGs in practical applications (Paulheim2016KnowledgeGR, ), where self-derived knowledge may not always be reliable and can be detrimental.
@@
Here, we adopt a matrix-projected distance to measure the distance between $\bar{\mathbf{e}}$ and $\mathbf{e}_{v}$, where $W^{1}_{E}\in\mathbb{R}^{d\times d}$ is the product of two learnable matrix $W_{E_{1}}\in\mathbb{R}^{d\times k},W_{E_{2}}\in\mathbb{R}^{k\times d}$ to enhance flexibility. $d$ is the dimension of $e$, and $k$ is a hyperparameter. We mapped $\bar{\mathbf{e}}$ and $\mathbf{e}_{v}$ into k-dimension and computed their distance via dot production. $\gamma$ is introduced to control the scaling, while $softmax(.)$ is utilized for normalization. Finally, the extracted sementic knowledge can be integrated as follows:
@@ 5. Experiments
  • RQ2: How does CSD compare with well-established distillation techniques frequently utilized in low-dimensional knowledge graph embedding?
@@
  • RQ5: What is the impact of these partial modules - iterative distillation, periodic snapshot distillation and confidence-aware semantic distillation - in enhancing the model’s efficiency?
@@ 5.1.1. Datasets
Link prediction experiments are performed on three standard benchmark datasets: YAGO3-10 (mahdisoltani2013yago3, ), FB15K-237 (FB15k237, ), and WN18RR (ConvE, ). These datasets have been extensively utilized in knowledge graph embedding studies. Note that datasets FB15K-237 and WN18RR are subsets of FB15k (TransE, ) and WN18 (TransE, ), respectively. FB15k and WN18 exhibit test data leakage, where numerous inverse relations from the testing sets can be conveniently discovered in the training sets (toutanova2015representing, ). FB15K-237 and WN18RR datasets were generated by eliminating inverse relations to rectify this problem. We therefore do not test our method on FB15k and WN18 datasets. Detailed statistics of these datasets are provided in Table 1.
@@
Table 2. Comparison results of link prediction on YAGO3-10, WN18RR and FB15K-237 datasets with 25 dimensions. The $Gain$ is the improvement of the CSD ratio relative to the baseline. The - indicates that the method cannot be applied in the corresponding model in our hardware due to out of memory.

### v3

@@ Abstract.
Knowledge Graph Embedding (KGE), which projects entities and relations into continuous vector spaces, has garnered significant attention. Although high-dimensional KGE methods offer better performance, they come at the expense of significant computation and memory overheads. Decreasing embedding dimensions significantly deteriorates model performance. While several recent efforts utilize knowledge distillation or non-Euclidean representation learning to augment the effectiveness of low-dimensional KGE, they either necessitate a pre-trained high-dimensional teacher model or involve complex non-Euclidean operations, thereby incurring considerable additional computational costs. To address this, this work proposes Confidence-aware Self-Knowledge Distillation (CSD) that learns from the model itself to enhance KGE in a low-dimensional space. Specifically, CSD extracts knowledge from embeddings in previous iterations, which would be utilized to supervise the learning of the model in the next iterations. Moreover, a specific semantic module is developed to filter reliable knowledge by estimating the confidence of previously learned embeddings. This straightforward strategy bypasses the need for time-consuming pre-training of teacher models and can be integrated into various KGE methods to improve their performance. Our comprehensive experiments on six KGE backbones and four datasets underscore the effectiveness of the proposed CSD.
Knowledge Graph Embedding, Knowledge Distillation, Confidence, Self-distillation
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679683isbn: 979-8-4007-0436-9/24/10ccs: Information systems Data mining
@@ 1. Introduction
Knowledge Graphs (KGs), which encapsulate real-world facts in a graph structure, have found extensive applications across various domains, ranging from event prediction (EP1, ; EP2, ), question answering (QE1, ; QE2, ), to recommender systems (RS1, ; RS2, ; cui2024distillation, ). As a foundational task in this area, Knowledge Graph Embedding (KGE) (TransE, ; wu2021disenkgat, ) aims to represent entities and relations in KGs as continuous vectors. With the increasing scale of KGs in practice, the necessity for both accuracy and efficiency in KGE becomes increasingly critical. While recent years have witnessed a proliferation of sophisticated KGE methods, most successes hinge on high-dimensional embeddings, which inevitably incur increased memory and computational overhead. For instance, as demonstrated in Figure 1, representative KGE methods such as LowFER (LowFER, ) and DistMult (DistMult, ) exhibit a significant performance decline (about 96%) in their low-dimensional versions (e.g., 25 dimensions) compared to their high-dimensional (e.g., 250 dimensions) counterparts, despite the low-dimensional models have 6-11 times fewer embedding layer parameters and offering 3-5 times improved inference latency.
To enhance the effectiveness of low-dimensional KGE, several recent efforts have emerged, which can be categorized into two main approaches: 1) leveraging knowledge distillation to empower a low-dimensional model with knowledge from high-dimensional teacher models (chen2023unbiased, ; liu2023iterde, ; MulDE, ; DualDE, ); and 2) embedding KGs in the non-Euclidean spaces (e.g., hyperbolic or mixed-curvature spaces) to better capture the complex patterns of KGs (cao2022geometry, ). However, these strategies still face limitations: knowledge distillation requires pre-training a high-dimensional KGE model, which still demands extensive computational resources and memory; the non-Euclidean methods involve complex operations that incur additional computational costs and lack flexibility to be applied in various KGE models (yao2023analogical, ). Given these limitations, an intriguing research question arises: How can we achieve effective low-dimensional KGE without incurring significant additional training costs?
@@
To this end, this work proposes to leverage self-knowledge distillation (SKD) for both efficient and effective knowledge graph embedding with low dimensions. Unlike traditional knowledge distillation (KD), which requires a pre-trained teacher model, SKD learns from the model itself through introspection (selfKD1, ; selfKD2, ; snapshot, ; yun2020regularizing, ). Specifically, SKD utilizes information from past epochs to extract valuable knowledge to guide learning in subsequent epochs. SKD has been proven to enhance model accuracy and robustness while avoiding the substantial training overhead, and has been successfully applied in fields such as computer vision and natural language processing. However, adapting sophisticated SKD strategies to KGE presents unique challenges:
@@
  • We propose an innovative iterative self-knowledge distillation method tailored for knowledge graph embedding that employs a specific semantic extraction module to distill reliable embedding knowledge from previous epochs to guide model learning.
@@ 2.1. Knowledge Graph Embedding
Recently, there has been intensive research on a neural network model called ConvE (ConvE, ), which aims to capture the semantic interactions between entities and relations in knowledge graphs. Building upon this, ConvR (ConvR, ) improves ConvE by replacing global convolution filters with relation-based filters. These filters are adaptively generated from relation embeddings, allowing ConvR to capture entity-relation interactions in the knowledge graph. Another model called HypER (HypER, ) further extends ConvE by using a hypernetwork to generate relation-specific convolution filters through tensor factorization. This extension facilitates feature interactions between entities and relations in knowledge graphs. Additionally, AcrE (AcrE, ) and InteractE (InteractE, ) introduce various convolution mechanisms, such as atrous and circular convolution, to enhance entity and relation semantic interactions. JointE (JointE, ) takes this further by jointly utilizing 1D and 2D convolutions to capture translation properties in KGs, thereby increasing interactions between entities and relations. M-DCN (M-DCN, ) introduces a dynamic and adaptive generation of multi-scale filters for convolutions. It captures feature interactions and improves model expressiveness in knowledge graphs.
@@ 2.2. Knowledge Distillation
To get both performance and inference speed, knowledge distillation technology has been proposed. The initial knowledge distillation approach focused on utilizing the teacher model’s prediction probabilities as soft a targets for training the student model (firstKD, ; wang2022semckd, ). FitNet (FitNet, ) advanced the process by using the intermediate features of the teacher model as training hints for the student model. Building on FitNet, SemCKD (SemCKD, ) introduced semantic calibration, which employed an attention mechanism to obtain adaptive intermediate features for student training. Additionally, SimKD (SimKD, ) achieved competitive performance for the student model simply by reusing the discriminant classifier from the teacher model.
The pioneer in knowledge graph embedding, KD-MKB (KD-MKB, ), first utilized knowledge distillation for knowledge graph embedding by establishing a cooperative learning framework between a teacher model and a student model. Following this, MulDE (MulDE, ) and DualDE (DualDE, ) are proposed. All these papers utilized high-dimensional, powerful teacher models to facilitate the training of student models for improved performance. However, a high-capacity teacher model is not available in many cases. To address this limitation, we shifting our focus towards self-knowledge distillation strategies(selfKD1, ; selfKD2, ; furlanello2018born, ; yun2020regularizing, ). However, existing self-distillation methods predominantly cater to the field of computer vision and yet perform sub-optimally within the field of knowledge graph embedding, where they manage input data that generally assume independence from each other. However, in the Knowledge Graph Embedding (KGE) domain, the data that we input are interconnected. The current self-distillation algorithms fail to leverage this aspect of information. Consequently, to harness this component of data, we propose the Improved Self-Distillation (CSD) method.
@@ 3.1. Knowledge Graph Embedding
A knowledge graph (YAGO, ), denoted as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{T}\}$, encompasses substantial quantities of entities and relations. Herein, $\mathcal{V}$, $\mathcal{R}$ and $\mathcal{T}$ signify the sets of entities, relations, and triplets respectively. Each fact $t\in\mathcal{T}$, is presented as a triple $(h,r,t)\in\mathcal{V}\times\mathcal{R}\times\mathcal{V}$, describing there is a relationship $r\in\mathcal{R}$ from the head entity $h\in\mathcal{V}$ to the tail entity $t\in\mathcal{V}$. The target of Knowledge graph embedding (KGE) is to project each entity $v\in\mathcal{V}$ and relation $r\in\mathcal{R}$ into corresponding d-dimensional representations $\mathbf{e}_{v},\mathbf{f}_{r}$. For convenience, we collect the embeddings of all entities and relations as matrices $\mathbf{E}$ and $\mathbf{F}$, respectively. These embeddings are subsequently utilized to facilitate downstream tasks.
To derive embeddings from Knowledge Graphs (KGs), existing Knowledge Graph Embedding (KGE) methods (TuckER, ; DistMult, ; ComplEx, ) typically formulate the task as a specific link prediction problem. They establish a scoring function $S_{hrt}=g(\mathbf{e}_{h},\mathbf{f}_{r},\mathbf{e}_{t})$ based on embeddings to estimate the validity of a triplet. For instance, TransE (TransE, ) adopts $g(\mathbf{e}_{h},\mathbf{f}_{r},\mathbf{e}_{t})=-\|\mathbf{e}_{h}+\mathbf{f}_{r}-\mathbf{e}_{t}\|_{1/2}$. Typically, the BCE loss can be used to optimize the KGE models as follows:
@@ 3.2. Self-Knowledge Distillation
(3) $D_{H}(a,b)=\begin{cases}\frac{1}{2}*(a-b)^{2},&|a-b|\leq 1\\ |a-b|-\frac{1}{2},&|a-b|>1.\end{cases}$
SKD employs a previous snapshot of the model as the supervises the learning of the current model.
@@ 4. Methodology
In this section, we thoroughly discuss the proposed Confidence-aware Self-Semantic Distillation (CSD), which is tailored for low-dimensional KGE.
As shown in Fig. 2 (a), CSD follow sthe conventional architecture of SKD, where a KGE model alternates between the roles of teacher and student during the training process. The knowledge of the previous iteration would be utilized to supervise the learning of the current model. However, given the specific challenges in KGE, we make specific designs to adapt SKD to this task.
Diverge from utilizing soft labels in the classification task, we distill knowledge from the model embeddings, by taking model status in $l-1$-th iteration as a teacher in $l$-th iteration. A specific semantic extraction block is developed to extract reliable knowledge from model embeddings to supervise the model learning (discussed in Section 4.1). Semantic information is dynamically and adaptively generated by the semantic extraction block, showcasing dynamic and evolution properties throughout the training process. Throughout these iterations, we leverage the teacher model to transfer semantic information to the student model by optimizing the Huber Loss between their semantic information. This optimization facilitates the student model to progressively incorporate and accumulate semantic knowledge from the past model.
@@ 4.1. Semantic Extraction Block
The embedding matrices $E$ and $F$ effectively encapsulate the knowledge graph’s semantic attributes based on the entities and relations. However, we find that directly distilling the embedding matrices between the teacher and student models leads to poor performance gain. This is due to the complex structures and inherently noisy nature of KGs in practical applications (Paulheim2016KnowledgeGR, ), where self-derived knowledge may not always be reliable and can be detrimental.
@@
Here, we adopt a matrix-projected distance to measure the distance between $\bar{\mathbf{e}}$ and $\mathbf{e}_{v}$, where $W^{1}_{E}\in\mathbb{R}^{d\times d}$ is the product of two learnable matrices $W_{E_{1}}\in\mathbb{R}^{d\times k},W_{E_{2}}\in\mathbb{R}^{k\times d}$ to enhance flexibility. $d$ is the dimension of $e$, and $k$ is a hyperparameter. We mapped $\bar{\mathbf{e}}$ and $\mathbf{e}_{v}$ into k-dimension and computed their distance via dot production. $\gamma$ is introduced to control the scaling, while $softmax(.)$ is utilized for normalization. Finally, the extracted semantic knowledge can be integrated as follows:
@@ 5. Experiments
  • RQ2: How does CSD compare with well-established distillation techniques frequently utilized in knowledge graph embedding?
@@
  • RQ5: What is the impact of these partial modules - iterative distillation, periodic snapshot distillation, and confidence-aware semantic distillation - in enhancing the model’s efficiency?
@@ 5.1.1. Datasets
Comprehensive link prediction experiments are performed on three standard benchmark datasets: YAGO3-10 (mahdisoltani2013yago3, ), FB15K-237 (FB15k237, ), and WN18RR (ConvE, ). These datasets have been extensively utilized in knowledge graph embedding studies. Note that datasets FB15K-237 and WN18RR are subsets of FB15k (TransE, ) and WN18 (TransE, ), respectively. FB15k and WN18 exhibit test data leakage, where numerous inverse relations from the testing sets can be conveniently discovered in the training sets (toutanova2015representing, ). FB15K-237 and WN18RR datasets were generated by eliminating inverse relations to rectify this problem. We therefore do not test our method on FB15k and WN18 datasets. Detailed statistics of these datasets are provided in Table 1.
@@
Table 2. Comparison results of link prediction on YAGO3-10, WN18RR and FB15K-237 datasets with 25 dimensions. The $Gain$ is the improvement of the CSD ratio relative to the baseline. The Baseline refers to the results obtained from directly training the models without applying any distillation methods. The - indicates that the method cannot be applied in the corresponding model in our hardware due to being out of memory.
@@
5.1.2. Baselines
5.1.3. Metrics
5.1.4. Implementation Details
We implement all models with the proposed CSD strategy in PyTorch. All experiments are conducted on an NVIDIA GeForce GTX 4090 GPU(24GB). For general link prediction experiments, we utilize the Adam (Adam, ) optimizer. Batch normalization (BN, ) and dropout (dropout, ) are used to mitigate over-fitting. To investigate the model performance in the low-dimension space, we set the embedding dimensions to 25 and 100. Other detailed parameter settings are approximately identical to their original papers. All results of * are retrieved from the paper of DualDE (DualDE, ) or IterDE (liu2023iterde, ). We trained 500 epochs for each knowledge graph model. For the CSD strategy, we set $L$ to $125$, select $k=50,\gamma=0.25$ for 25 dimensions and $k=200,\gamma=0.1$ for 100 dimensions. We search the value of $\beta$ and $\lambda$ in the interval $(0,1)$. For every sub-training process, $\beta$ will be reset to the initial setting value and decay over training epochs to take advantage of the temporal attentiveness. The value of $\beta$ multiplied by $(1-ep/EP)$ at the beginning of $ep$-th epoch in this sub-training process, where $EP$ is the total number of epochs for each sub-training process.
5.2. Link Prediction Results (RQ1 and RQ2)
We perform link prediction experiments for all baselines combined with our method. The experimental results of 50 and 100 dimensions are shown in Tables 2 and 3. The best results are in bold. Overall, the baseline model combined with our method outperforms all original baselines on these datasets. Specifically, the LowFER model combined with CSD with 25 dimensions has significant performance gains on the WN18RR dataset with improvements of 279.4% (from 0.0160 to 0.0607), 292.2% (from 0.0282 to 0.1106) and 339.7% (from 0.0078 to 0.0343) in MRR, H@10 and H@1, respectively.
Comparing CSD with CS-KD and LWR. Our CSD surpasses both CS-KD and LWR across all backbones and datasets, thereby substantiating CSD’s efficacy. It merits mentioning that employing the comparative method, LWR, on the YAGO3-10 dataset proves challenging due to the necessity of recording the model’s output logits for each sample, resulting in considerable memory consumption. And the comparative methods CS-KD and LWR do not surpass the baseline performance. CS-KD is primarily designed for image classification tasks involving limited classes and encounters challenges in the link prediction task. The large number of entities in this task significantly exceeds the class quantity in image classification tasks, resulting in insufficient samples for CS-KD to acquire class-specific knowledge, thereby leading to performance degradation. LWR’s approach to directly distill knowledge from a weak prediction to the model proves effective in image classification tasks, owed primarily to the dark knowledge contained in the weak prediction distribution. Conversely, in the link prediction task, the dark knowledge of a weak prediction distribution declines due to the immense number of entities involved. Consequently, the detrimental impact of a weak prediction distribution outweighs its benefit, some erroneous prediction distribution for knowledge distillation may cause the student model to undergo performance degradation.
Comparing CSD with SD. Our proposed CSD outperforms SD on almost every backbone and dataset. SD, akin to our proposed CSD, distills knowledge of the teacher model obtained through the periodic learning rate strategy. Despite SD yielding satisfactory results on certain models, some instances revealed that the Mean Reciprocal Rank (MRR) performance was inferior to the Baseline. The inclusion of a semantic extraction block facilitates the student model in acquiring semantic knowledge based on the last iteration. This Additional advantage allows CSD to surpass SD in performance across the majority of models and datasets.